OpenStudy (anonymous):

medal for best answer Which is a solution of the equation? y = –3x + 5 (2, 1) (–1, –8) (3, –4) (–2, 0)

OpenStudy (anonymous):

Plug each in to find the answer,

OpenStudy (anonymous):

1=-3(2)+5 solve.

OpenStudy (anonymous):

-8=-3(-1)+5 solve.

OpenStudy (anonymous):

-4=-3(3)+5 solve.

OpenStudy (anonymous):

0=-3(2)+5 solve.

OpenStudy (anonymous):

1=-6+5 1=-1 NO. -8=-9+5 -8=-4 NO. -4=-9+5 -4=-4 Yes! 0=-6+5 0=-1 NO.